# Mechanisms and Robotics

### v1

@@ Abstract
A cuspidal robot can move from one inverse kinematics (IK) solution to another without crossing a singularity. Multiple industrial robots are cuspidal. They tend to have a beautiful mechanical design, but they pose path planning challenges. A task-space path may have a valid IK solution for each point along the path, but a continuous joint-space path may depend on the choice of the IK solution or even be infeasible. This paper presents new analysis, path planning, and optimization methods to enhance the utility of cuspidal robots. We first demonstrate an efficient method to identify cuspidal robots and show, for the first time, that the ABB GoFa and certain robots with three parallel joint axes are cuspidal. We then propose a new path planning method for cuspidal robots by finding all IK solutions for each point along a task-space path and constructing a graph to connect each vertex corresponding to an IK solution. Graph edges are weighted based on the optimization metric, such as minimizing joint velocity. The optimal feasible path is the shortest path in the graph. This method can find non-singular paths as well as smooth paths which pass through singularities. Finally, this path planning method is incorporated into a path optimization algorithm. Given a fixed workspace toolpath, we optimize the offset of the toolpath in the robot base frame while ensuring continuous joint motion. Code examples are available in a publicly accessible repository.
@@ 1 Introduction
The most common design for large industrial robot arms features a spherical wrist and two parallel axes. Examples include the ABB IRB 6640 [1] and the YASKAWA Motoman MA2010-A0 [2]. Cobots designed for collaborative tasks such as the UR5 [3] or Techman TM5-700 [4] feature three parallel axes plus other intersecting axes. In either case, a given end effector pose has up to eight inverse kinematics (IK) solutions which are found by solving a series of quadratic equations. The IK solutions can be intuitively labeled and are separated by singularities.
@@
Cuspidal robots can travel between IK solutions without encountering a singularity. They have IK that cannot be solved with quadratics and have certain constraints on kinematic parameters such as link lengths. For a long time, this property was thought to be impossible, and an incorrect proof of this impossibility was published [8], but this proof was later proven wrong [9]. Even today, robot cuspidality is rarely considered [7] but a robot with a randomly chosen geometry is almost surely cuspidal [10]. It appears cuspidal robots are being used without manufacturers or practitioners realizing they are cuspidal or fully appreciating the resulting path planning challenges.
The following example illustrates a case where cuspidality can cause path planning issues. Consider the case of robotic welding, which is an advertised use case for many industrial cuspidal robots. A robot operator may program a closed path for a welding operation, and there may be no apparent problem because the robot can perform the entire weld. However, since the robot is cuspidal, the robot can perform a nonsingular change of solution during the welding operation. This means when the robot executes the weld for a second time, the robot may get stuck at a singularity, or, if the first command in the program is a joint-motion command, the robot may produce unexpected large motion to return to the original configuration!
@@
Cuspidal robots highlight the importance of considering all IK solutions during path planning and demonstrate the need for IK solvers which can find all IK solutions. (In fact, finding all IK solutions is important for path planning even for noncuspidal robots as certain paths may be non-optimal or even infeasible depending on the choice of initial IK solution.) Many IK solvers are Jacobian-based iterative solvers, and these do not return all IK solutions.
IK solvers used by industrial robots which return multiple solutions may have problems. ROBOGUIDE, the software used by FANUC to control their CRX line of robots, only provides up to eight IK solutions even though the CRX has up to sixteen in general [10]. Furthermore, IK solutions in ROBOGUIDE are labeled using shoulder/elbow/wrist configuration parameters, but these descriptions may not correctly match the configuration of the arm and may not change when changing IK solutions. In fact, we can find end effector poses for the CRX with sixteen IK solutions but only two aspects [11]! In Section 2.2, we show very similar issues for the ABB GoFa (Fig. 1(c)) for the first time.
@@
The remainder of this paper is structured as follows. We provide definitions, give an overview of cuspidal path planning challenges, and discuss related literature in Section 2. We identify cuspidal robots in in Section 3. We develop and demonstrate path planning method in Section 4 and use it for path optimization in Section 5. We conclude in Section 6.
@@ 2 Background
Figure 2: Cuspidal 3R manipulator with an infeasible path (straight line in task space) and a nonsingular change of solution (straight line in joint space). (a) 3D view. Points with singularities are marked for the $\phi=0$, $\rho>0$ half-plane. (b) All solutions for joint angle 1 over the infeasible path. Line width corresponds to which aspect the IK solution is in. (c) One aspect in joint space bounded by singularities. Nonsingular change of solution is shown. (d) Forward kinematics mapping of one aspect to task space. Also shown are the infeasible path and the nonsingular change of solution.
@@ 2.1 Definitions
Points in the characteristic surface of an aspect (in joint space) correspond to end effector poses that have a singular solution in the same aspect. Noncuspidal robots have no characteristic surfaces.
@@
Parallel robots can also be cuspidal [7], but the roles of the forward and inverse kinematics mapping are reversed. Cuspidal parallel robots have closed active-joint-space paths with different beginning and ending end effector poses.
@@ 2.2 Cuspidal Robot Path Planning Challenge
In the product of exponentials convention (which is further described in [12]), $p_{01}$ is the position of the first link frame in the base frame, $p_{i-1,i}$ is the position of link frame $i$ in the $i-1$ frame, and $p_{i,T}$ is the position of the end effector task frame in the $i$ frame. $h_{i}$ is the joint axis directions for joint $i$ represented in the $i-1$ frame. $e_{x}$, $e_{y}$, and $e_{z}$ are the Cartesian unit vectors.
@@
We demonstrate cuspidal path planning issues for the ABB GoFa in RobotStudio and the FANUC CRX-10iA/L in ROBOGUIDE.
@@
We also raise the concern for the first time that RobotStudio does not find all IK solutions for the GoFa. The robot has up to 16 solutions, but only up to 8 solutions are returned. That the solutions are labeled from 0 to 7 is explicitly called out for the GoFa in the programming manual [16, Sec. 3.16].
@@ 2.3 Related Literature
Cuspidal manipulators have been known for some time. Although there is limited research in this field, there has been a recent resurgence in interest because cuspidal manipulators are becoming commercially widespread for the first time.
Most research discusses issues arising from manipulators being cuspidal, and many papers conclude cuspidal robots should be avoided. For example, [11] argued cuspidal robots should be avoided in collaborative applications, and [7] argued it is reasonable to avoid designing cuspidal robots.
@@ 2.3.1 Identification of Cuspidal Manipulators
If the robot can be solved without using polynomials of degree three or higher, then the robot is not cuspidal, or, equivalently, if the determinant factors into three terms, then the robot is noncuspidal. A list of noncuspidal 6R robots with this property is found in [11]. There are many kinematic simplifications that can result in a polynomial of degree two [17]. Two common cases for revolute manipulators to be solved with only quadratics is to have a spherical wrist or three parallel axes plus two other intersecting or parallel axes [12]. If a 6R robot has a spherical wrist, then it is cuspidal if and only if the 3R robot formed by the first three links is cuspidal.
A 3R robot is cuspidal if and only if it has a cusp point in the locus of singular points in the workspace, which is equivalently a triple point with exactly three solutions. Necessary and sufficient kinematic conditions for cuspidality have been found for 3R orthogonal manipulators [7]. It has only been recently proven that a general 3R manipulator is cuspidal if and only if it has a cusp point [18]. It is unknown if the same condition holds for 6R manipulators, and this condition is false for parallel manipulators. It is unfortunate that the name “cuspidal" appears to be a misnomer as cusp points are not necessary in general for a manipulator to be cuspidal.
@@
The identification of 6R manipulators has been reduced to the path planning problem of finding a nonsingular path between IK solutions, which we discuss further in Section 2.3.2. [15] identified cuspidal robots by finding a linear joint-space MoveJ path between IK solutions and then performing path optimization to avoid passing through a singularity twice. [10] identified robots by solving the optimal path planning problem with a multiple shooting approach. Finding a nonsingular change of IK solution proves a robot is cuspidal. Not finding such a path means a robot is likely noncuspidal, but this is not proof.
@@ 2.3.2 Path Planning for Cuspidal Manipulators
A different path planning problem is finding a feasible joint-space path for a given task-space path. This can be for open paths or closed paths. To the best of our knowledge, there has only been one path planning algorithm proposed for cuspidal robots following a prescribed task-space path presented in [10, 11] (the algorithms, presented as flow charts, are slightly different between [10] and [11] in the case that the path doesn’t cross a characteristic surface.) The algorithm is:
@@
    Pick one IK solution along the path with the smallest number of IK solutions
@@
    If the path does not cross a characteristic surface, check if the path is continuous
@@ 2.3.3 Path Optimization for Cuspidal Manipulators
Although there exist algorithms for noncuspidal robots for optimizing the rigid-body transformation of a task-space path, no optimization algorithms have been proposed specifically for cuspidal robots before now. Path planning algorithms for noncuspidal robots may not work correctly for cuspidal robots because they may not consider all IK solutions, may not check for changes of solutions for closed task-space paths, or may check for connectivity based on aspect or the presence of IK solutions throughout the path.
Proposed path-planning algorithms for non-cuspidal robots are often related to machining tasks such as drilling or milling where robot stiffness must be optimized. These tasks are achieved using robots with spherical wrists [23, 24, 25, 26, 27] or with a UR robot [28, 29]. Optimization was done for a Stewart platform in [30]. Some literature resolves the redundant degree of freedom about a spinning tool [25, 26]. In [27], the relative path of two robots was optimized, which required resolving six redundant degrees of freedom. Many algorithms use optimization strategies like particle swarm optimization (PSO) [24, 26] or genetic algorithms [30]. Objective functions included optimizing joint velocities [29, 23], cycle time [27], joint torque and manipulability [28], stiffness  [26, 30], contour accuracy (by considering cutting forces and stiffness) [24, 25], and power consumed [30]. Tool offset was optimized in addition to base position in [23]. Furthermore, [23] proposed a path planner similar to ours (it is implicitly a graph search) that can go through singularities. Rather than optimizing over six degrees of freedom, [29] optimized only over the three degrees of freedom in the placement of a workpiece on a table. Euler angles were used for parameterization in [30, 26], but the singularity problem was not addressed.
@@ 3.2 Solution Method
    Pick a random pose.
@@
In [10], cuspidality was determined by finding all IK solutions using the Husty-Pfurner algorithm, which took on average 10 ms per end effector pose, and then connecting them using a multiple shooting approach, which took 8.52 sec. In contrast, IK for a single end effector pose using IK-Geo takes under 1 ms for robots with at least one pair of intersecting or parallel axes.

### v2

@@ Abstract
A cuspidal robot can move from one inverse kinematics (IK) solution to another without crossing a singularity. Multiple industrial robots are cuspidal. They tend to have a beautiful mechanical design, but they pose path planning challenges. A task-space path may have a valid IK solution for each point along the path, but a continuous joint-space path may depend on the choice of the IK solution or even be infeasible. This paper presents new analysis, path planning, and optimization methods to enhance the utility of cuspidal robots. We first demonstrate an efficient method to identify cuspidal robots and show, for the first time, that the ABB GoFa and certain robots with three parallel joint axes are cuspidal. We then propose a new path planning method for cuspidal robots by finding all IK solutions for each point along a task-space path and constructing a graph to connect each vertex corresponding to an IK solution. Graph edges have a weight based on the optimization metric, such as minimizing joint velocity. The optimal feasible path is the shortest path in the graph. This method can find non-singular paths as well as smooth paths which pass through singularities. Finally, we incorporate this path planning method into a path optimization algorithm. Given a fixed workspace toolpath, we optimize the offset of the toolpath in the robot base frame while ensuring continuous joint motion. Code examples are available in a publicly accessible repository.
@@ 1 Introduction
The most common design for large industrial robot arms features a spherical wrist and two parallel axes. Examples include the ABB IRB 6640 [1] and the YASKAWA Motoman MA2010-A0 [2]. Cobots designed for collaborative tasks such as the UR5 [3] or Techman TM5-700 [4] feature three parallel axes plus other intersecting axes. In either case, a given end effector pose has up to eight inverse kinematics (IK) solutions which correspond to the solutions of a series of quadratic equations. The IK solutions have intuitive labels and are separated by singularities.
Cuspidal robots can travel between IK solutions without encountering a singularity. They have IK that cannot be solved with quadratics and have certain constraints on kinematic parameters such as link lengths. For a long time, this property was thought to be impossible. An incorrect proof of this impossibility was published [8], but this proof was later proven wrong [9]. Even today, robot cuspidality is rarely considered [7] but a robot with a randomly chosen geometry is almost surely cuspidal [10]. It appears manufacturers and practitioners are building and using cuspidal robots without realizing they are cuspidal or fully appreciating the resulting path planning challenges.
Although there is limited research in this field, there has been a recent resurgence in interest because cuspidal manipulators are becoming commercially widespread for the first time. Most research discusses issues arising from manipulators being cuspidal, and many papers conclude cuspidal robots should be avoided. For example, [11] argued cuspidal robots should be avoided in collaborative applications, and [7] argued it is reasonable to avoid designing cuspidal robots.
@@
The following example illustrates a case where cuspidality can cause path planning issues. Consider the case of robotic welding, which is an advertised use case for many industrial cuspidal robots. A robot operator may program a closed path for a welding operation, and there may be no apparent problem because the robot can perform the entire weld. However, since the robot is cuspidal, the robot can perform a nonsingular change of solution during the welding operation. This means when the robot executes the weld for a second time, the robot may get stuck at a singularity. Moreover, if the first command in the program is a joint-motion command, the robot may produce unexpected large motion to return to the original configuration!
@@
Cuspidal robots highlight the importance of considering all IK solutions during path planning and demonstrate the need for IK solvers which can find all IK solutions. Missing IK solutions is problematic because feasible paths may be missed, resulting in a failure to connect two points or picking a suboptimal path. (In fact, finding all IK solutions is important for path planning even for noncuspidal robots as certain paths may be non-optimal or even infeasible depending on the choice of initial IK solution.) Many IK solvers are Jacobian-based iterative solvers, and these do not return all IK solutions.
IK solvers used by industrial robots which return multiple solutions may have problems. ROBOGUIDE, the software used by FANUC to control their CRX line of robots, only provides up to eight IK solutions even though the CRX has up to sixteen in general [10]. Furthermore, ROBOGUIDE labels IK solutions using shoulder/elbow/wrist configuration parameters, but these descriptions may not correctly match the configuration of the arm and may not change when changing IK solutions. In fact, we can find end effector poses for the CRX with sixteen IK solutions but only two aspects [11]. In Section 2.2, we show very similar issues for the ABB GoFa (Fig. 1(c)) for the first time.
@@
By leveraging the efficient and complete IK solutions from IK-Geo, the goal of this paper is to identify cuspidal robots and perform path planning and path optimization for these robots. Path planning algorithms for noncuspidal robots may not work correctly for cuspidal robots because they may not consider all IK solutions, may not check for changes of solutions for closed task-space paths, or may check for connectivity based on aspect or the presence of IK solutions throughout the path. Although there exist algorithms for noncuspidal robots for optimizing the rigid-body transformation of a task-space path, no optimization algorithms have been proposed specifically for cuspidal robots before now.
@@
The remainder of this paper is structured as follows. We explain our notation for robot kinematics, provide definitions, give an overview of cuspidal path planning challenges, and discuss related literature in Section 2. We identify cuspidal robots in in Section 3. We develop and demonstrate path planning method in Section 4 and use it for path optimization in Section 5. We conclude in Section 6.
@@ 2 Background
Figure 2: Cuspidal 3R manipulator with an infeasible path (straight line in task space) and a nonsingular change of solution (straight line in joint space). (a) 3D view using cylindrical coordinates $(\rho,\phi,z)$. Points with singularities are marked for the $\phi=0$, $\rho>0$ half-plane. (b) All solutions for joint angle 1 over the infeasible path. Line width corresponds to which aspect the IK solution is in. (c) One aspect in joint space bounded by singularities. Nonsingular change of solution is shown. (d) Forward kinematics mapping of one aspect to task space. Also shown are the infeasible path and the nonsingular change of solution.
The configuration of a serial revolute robot manipulator is determined by the joint angle vector
$q=[q_{1}\ q_{2}\ \dots\ q_{6}]^{T}\in\mathbb{R}^{6},$ (1)
where in this case the robot has six degrees of freedom. A 6-DOF revolute manipulator is denoted 6R, and a 3-DOF revolute manipulator is denoted 3R. We will consider nonredundant manipulators in this paper, meaning the joint space and task space have the same dimension.
If we consider joint angles $q_{i}$ and $q_{i}+2\pi$ equivalent, such as when there are no joint limits, then we can model the joint space as a torus, that is, $q\in S^{6}$ and $q_{i}\in[-\pi,\pi]$.
Forward kinematics $f_{0T}$ is the mapping from joint space to task space:
$\chi_{0T}=(R_{0T},p_{0T})=f_{0T}(q),$ (2)
where $\chi_{0T}\in\mathrm{SE}(3)$ is the pose of the task (end effector) frame in the robot base frame. This consists of the end effector orientation $R_{0T}\in\mathrm{SO}(3)$ and Cartesian end effector position $p_{0T}\in\mathbb{R}^{3}$. For a 3-DOF robot, the end effector pose only consists of the position. In addition to using Cartesian coordinates $(x,y,z)$, we may also represent the end effector position using cylindrical coordinates $(\rho,\phi,z)$, where $\rho=(x^{2}+y^{2})^{1/2}$ and $\phi=\mathrm{ATAN2}(y,x)$.
Inverse kinematics (IK) is the mapping from a given end effector pose to its preimage, which is the set of all joint angle vectors achieving that end effector pose:
$f_{0T}^{-1}(\chi_{0T})=\{q\in\mathbb{R}^{6}\mid f_{0T}(q)=\chi_{0T}\}.$ (3)
There may be multiple IK solutions or even a continuum of solutions for a given end effector pose.
The kinematic Jacobian $J$, which is a function of $q$, defines the linear mapping from joint velocity to task velocity:
$\nu_{0T}=[\mathscr{v}_{0T}\ \omega_{0T}]^{T}=J(q)\dot{q},$ (4)
where $\nu_{0T}\in\mathbb{R}^{6}\cong\mathfrak{se}(3)$ is the end effector spatial velocity composed of the linear velocity $\mathscr{v}_{0T}\in\mathbb{R}^{3}$ and the angular velocity $\omega_{0T}\in\mathbb{R}^{3}$. For a 3-DOF robot, the end effector velocity is just the linear velocity.
A kinematic singularity occurs when $J(q)$ loses rank, which occurs when $\det(J)=0$. At a singular configuration, there is some spacial velocity direction which cannot be achieved, and there is some direction in joint space which results in zero spatial velocity (perhaps only instantaneously). This internal robot motion without moving the end effector is called self-motion.
A task-space path $\underline{\chi}_{0T}(\lambda)$ is a mapping from a path variable $\lambda\in\mathbb{R}$ to an end effector pose. A joint-space path $\underline{q}(\lambda)$ is defined similarly. Whereas paths are indexed by the path variable $\lambda$, trajectories are indexed by time $t$.
@@ 2.1 Definitions
Points in the characteristic surface of an aspect (in joint space) correspond to end effector poses that have a singular solution bounding same aspect. Noncuspidal robots have no characteristic surfaces.
@@
Parallel robots can also be cuspidal [7, 14, 15, 16, 17], but the roles of the forward and inverse kinematics mapping are reversed. Cuspidal parallel robots have closed active-joint-space paths with different beginning and ending end effector poses.
@@ 2.2 Cuspidal Robot Path Planning Challenge
In the product of exponentials convention (which is further described in [12]), $p_{01}$ is the position of the first link frame in the base frame, $p_{i-1,i}$ is the position of link frame $i$ in the $i-1$ frame, and $p_{i,T}$ is the position of the end effector task frame in the $i$ frame. The joint axis direction for joint $i$ represented in the $i-1$ frame is denoted by $h_{i}$. The Cartesian unit vectors are denoted by $e_{x}$, $e_{y}$, and $e_{z}$.
@@
We demonstrate cuspidal path planning issues for the ABB GoFa in RobotStudio and the FANUC CRX-10iA/L in ROBOGUIDE. These are two newer robots from large manufacturers. This means the simulators are highly realistic, as they use virtual controllers which match the behavior of the real robot. This also means there is a practical application since these robots are currently in production. As neither robot has three intersecting or parallel joint axes, their IK is not solved in closed form, making them both interesting and challenging examples.
@@
We also raise the concern for the first time that RobotStudio does not find all IK solutions for the GoFa. The robot has up to 16 solutions, but RobotStudio returns only up to 8 solutions. That the solutions are labeled from 0 to 7 is explicitly called out for the GoFa in the programming manual [20, Sec. 3.16].
@@ 2.3.1 Identification of Cuspidal Manipulators
If the robot can be solved without using polynomials of degree three or higher, then the robot is not cuspidal, or, equivalently, if the determinant $\det(J)$ factors into three terms, then the robot is noncuspidal. A list of noncuspidal 6R robots with this property is found in [11]. Many kinematic simplifications result in a polynomial of degree two [21]. Two common cases for revolute manipulators to be solved with only quadratics is to have a spherical wrist or three parallel axes plus two other intersecting or parallel axes [12]. If a 6R robot has a spherical wrist, then it is cuspidal if and only if the 3R robot formed by the first three links is cuspidal.
A 3R robot is cuspidal if and only if it has a cusp point in the locus of singular points in the workspace, which is equivalently a triple point with exactly three solutions. The IK problem for a 3R manipulator (Subproblem 5 in [12]) can be solved as the intersection of two ellipses, and the singularities correspond to tangencies between ellipses [22]. A cusp point corresponds to osculating contact between two ellipses. Necessary and sufficient kinematic conditions for cuspidality have been found for 3R orthogonal manipulators [7]. Researchers have only recently proven that a general 3R manipulator is cuspidal if and only if it has a cusp point [23].
It is unknown if 6R manipulators are cuspidal if and only if they have a cusp point, and this condition is false for parallel manipulators. In parallel manipulators, a nonsingular change of solution does not require encircling a cusp point and can instead occur due to encircling an $\alpha$-curve [16, 22]. Unfortunately, the name “cuspidal" appears to be a misnomer as cusp points are not necessary in general for a manipulator to be cuspidal.
Even 2-DOF manipulators can be cuspidal. In particular, planar 2-X manipulators can be cuspidal [24]. These are tensegrity structures with two X-mechanisms, which have variable instantaneous centers of rotation, connected in series.
@@
The problem of identifying if a 6R manipulator is cuspidal has been reduced to a path planning problem: A robot is cuspidal if a path planner can find a nonsingular path between two different IK solutions of the same end effector pose. We discuss the path planning problem further in Section 2.3.2. One proposed identification method [19] used linear joint-space MoveJ paths as the initial guess for a path optimization algorithm that adjusted the path to avoid singularities. A subsequent study [10] identified cuspidal robots by formulating the path planning problem as an optimization problem and solving it with a multiple shooting approach. If a nonsingular path between IK solutions is found, this proves the robot is cuspidal. Conversely, failing to find such a path suggests the robot is likely noncuspidal, though this does not constitute a formal proof.
@@ 2.3.2 Path Planning for Cuspidal Manipulators
A different path planning problem is finding a feasible joint-space path for a given task-space path. This can be for open paths or closed paths. This is the problem we will solve in Section 4. To the best of our knowledge, there has only been one path planning algorithm proposed for cuspidal robots following a prescribed task-space path presented in [10, 11] (the algorithms, presented as flow charts, are slightly different between [10] and [11] in the case that the path doesn’t cross a characteristic surface.) The algorithm is:
@@
    Pick one IK solution along the path with the smallest number of IK solutions.
@@
    If the path does not cross a characteristic surface, check if the path is continuous.
@@ 2.3.3 Path Optimization for Cuspidal Manipulators
Proposed path-planning algorithms for non-cuspidal robots are often related to machining tasks such as drilling or milling where robot stiffness must be optimized. These tasks are commonly achieved using robots with spherical wrists [29, 30, 31, 32, 33] or with a UR robot [34, 35]. Optimization was done for a Stewart platform in [36]. Some literature resolves the redundant degree of freedom about a spinning tool [31, 32]. In [33], the relative path of two robots was optimized, which required resolving six redundant degrees of freedom. Many algorithms used optimization strategies like particle swarm optimization (PSO) [30, 32] or genetic algorithms [36]. Objective functions included optimizing joint velocities [35, 29], cycle time [33], joint torque and manipulability [34], stiffness  [32, 36], contour accuracy (by considering cutting forces and stiffness) [30, 31], and power consumed [36]. Tool offset was optimized in addition to base position in [29]. Furthermore, [29] proposed a path planner similar to ours (it is implicitly a graph search) that can go through singularities. Rather than optimizing over six degrees of freedom, [35] optimized only over the three degrees of freedom in the placement of a workpiece on a table. Euler angles were used for parameterization in [36, 32], but the representation singularity problem was not addressed.
@@ 3.2 Solution Method
    Pick a random end effector pose. This can be done by finding the forward kinematics of a random joint angle vector.
@@
In [10], the authors determined if a robot was cuspidal by finding all IK solutions using the Husty-Pfurner algorithm and then connecting them using a multiple shooting approach. Finding all IK solutions took on average 10 ms per end effector pose, and the multiple shooting approach took 8.52 sec. In contrast, IK for a single end effector pose using IK-Geo takes under 1 ms for robots with at least one pair of intersecting or parallel axes.